A force is applied to a 4kg object so the it produces an acceleration of 1.1m/s^2. However, there is friction on the surface it slides on; the coefficient of kinetic friction on the surface is .4, what is the magnitude of the force applied?

A. 30.08N

B. 5.08N

C. 10.08N

D. 20.08N

E. None of the above
Answer #1

net force = mass * acceleration

= ma

= 4*1.1 = 4.4 N

friction force= uN = umg = 0.4*4*9.8 = 15.68 N

Let F be the magnitude of force applied

F - Ff = net force

F -15.68 = 4.4

F = 4.4 + 15.68 = 20.08 N

so option D) 20.08 N is the correct answer